OpenStudy (brucebaner):

Manuel builds and sells wooden crates. He made a table to show how much his crates cost. Which equation represents the situation? Manuel’s Crates Crates 2 4 6 8 Cost 25 45 65 85 A. y – 2 = −10(x – 25) B. y – 4 = 10(x – 45) C. y – 65 = −10(x – 6) D. y – 85 = 10(x – 8)

OpenStudy (brucebaner):

can you help me @Michele_Laino

OpenStudy (michele_laino):

you have to construct this relationship: y=kx+h. inserting first point, we can write: 25=2k+h. inserting the second couple of your data, we can write: 45=4k+h. now we solve the first equation in order to find h: h=25-2k

OpenStudy (michele_laino):

then we substitute h=25-2k in the second equation: 45=4k+25-2k from which: 2k=20 then k=10. and h=25-2*10=5 so your function is: y=10x+5

OpenStudy (brucebaner):

so i don't understand can you plz explain again

OpenStudy (michele_laino):

ok! you have to determine both coefficient k, h in the relationship y=kx+h

OpenStudy (michele_laino):

so we use the first couple of your numerical dat, namely 2, 25. I supposed that x= number of crates, and y= cost of crates. substituting x=2 and y=25, we have: 25=2k+h

OpenStudy (michele_laino):

now we do the same with the second couple of your numerical data, namely x=4, y=45, so we can write; 45=4k+h. We have the subsequent system: \[25=2k+h, 45=4k+h\] using the nsubstitution method, we find: k=10, and h=5, so your function is: y=10x+5, namely D.

OpenStudy (brucebaner):

so its D @Michele_Laino

OpenStudy (michele_laino):

yes it is!

OpenStudy (brucebaner):

thx so so much i really appreciate it so much @Michele_Laino thx thx

OpenStudy (michele_laino):

thank you! @brucebaner

OpenStudy (brucebaner):

and thx again